What is iota^iota . I really needed to know this so my friend told me that ask the iitians . So i am asking an iitian , hopefully you will answer my question ASAP thanks

Grade:11

3 Answers

SAHIL
7 years ago
Iota is a complex number with the value of root under -1 , so iota to the power iota is , root under -1 to the power root under -1.
THANKS
POST IF ANY QUERY OR DOUBT
 
 
krunal
7 years ago
actually the ans is (-pie/2).
here, lets assume that A=i^i. 
taking log on both side, we can write> logA =ilogi
logA=ilog(1i)......
logA=i[ log1+logi]...........by logarithm property
logA=I[0+i(PIE/2)]......(ARG OF i=pie by 2)
on soloving we get logA=(-PIE/2)
THEREFORE A=e^(-pie/2)
Umakant biswal
7 years ago
@ lvanish 
iota is besically a greek letter which is used to define the imaginary part of a complex number 
lets say we have an equation x^2 +1 =0 
the value of x will be the square root of -1 which is fundamentally not possible . 
to solve such equation u need the square root of some negative number ,. 
the concept of complex number are introduced for that . 
iota is used to denote the square root of -1 . 
some books u will find that iota is a number whoose square is -1 ..